OpenStudy (anonymous):

The displacement (in meters) of a particle moving in a straight line is given by s = t^2–9t + 17. Find the instantaneous velocity when t = 4?

OpenStudy (anonymous):

differentiate it.. and put t=4

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

ds/dt=2t-9 now,when t=4, 2*4-9=ds/dt -1=velocity